OpenStudy (aaronandyson):

the difference between the two acute angles of right angled right triangle is 2pi/5 radians.Find them in radians.

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (anonymouse2015):

find in radians or degrees?

OpenStudy (aaronandyson):

radians

OpenStudy (roxy.girl):

@Qwertty123

OpenStudy (anonymouse2015):

Um I'm not sure sorry, I can only work out degrees sorry :(

OpenStudy (qwertty123):

The sum is 90 degrees and their difference is 2 pi/5 radians or 72 degrees. All you did was convert the difference to degrees. A + B = 90 A - B = 72 2A = 162 A = 81 B = 90 - 81 =?

OpenStudy (aaronandyson):

I didn't get you.

OpenStudy (qwertty123):

what is 90-81?

OpenStudy (aaronandyson):

The right angle is 90 degree so shouldn't i convert it to radians?

OpenStudy (aaronandyson):

90-81 = 9

OpenStudy (qwertty123):

whoops yes XD

OpenStudy (aaronandyson):

how do I do that?

OpenStudy (qwertty123):

oh okay Radian measure of a central angle is defined as the ratio of the arc length to the length of the radius. This definition leads to the arc length formula: . Radians and degrees are connected by the relationship . If you wish to convert from degrees to radians, multiply the degree measure by...~ http://www.montereyinstitute.org/courses/DevelopmentalMath/TEXTGROUP-1-19_RESOURCE/U19_L2_T1_text_container.html

OpenStudy (aaronandyson):

???

OpenStudy (aaronandyson):

Is the final answer 90 and 10 rads?

OpenStudy (peachpi):

no. let the two angles be a and b. All the angles in a triangle add up to π radians, so the two acute angles add up to π/2. So one equation you can write is \[a+b=\frac{ \pi }{ 2 }\] You're also given the difference, so another equation is\[a-b=\frac{ 2\pi }{ 5 }\] Solve the system of equations for a and b

OpenStudy (peachpi):

start by adding the equations together to eliminate b.

OpenStudy (aaronandyson):

b gets cancelled and a is 81?

OpenStudy (peachpi):

yes if you follow @Qwertty123's work, but there's no reason to convert anything to degrees since your answer has to be in radians. You'd have to convert the 81 to radians

OpenStudy (aaronandyson):

thats 90 rad?

OpenStudy (qwertty123):

Very well done! :3